Algebra: How much of a 14% acid solution should be added to 90 oz. of a 47% solution to get one of 29%?




Answer by mananth(16946)

-------------- percent ---------------- Amount
Acid Solution I 47 ---------------- 90 Oz
Acid Solution II 14 ---------------- x Oz
Mixture 29 ---------------- 90 + x Oz

47*90+14x=29(90 +x)
4230+14 x=2610 +29x
14x-29x =-4230+ 2610
-15x=-1620
/-15
x=108 Oz Acid Solution II

m.ananth@hotmail.ca